OpenStudy (anonymous):

log a=log (4a-9)??????

OpenStudy (anonymous):

a=3

OpenStudy (anonymous):

log a=log (4a-9) log a-log (4a-9)=0 log (a/4a-9)=0 a/(4a-9)=1 a=4a-9 4a-a=9 3a=9 a=3